OpenStudy (anonymous):

On the plane xy, there are 4 points: A(a, b), B(-1,0), C(2,1), & D(0,2). If ∠ABC = 90' & point D is on the side AC, then a & b = ?

OpenStudy (anonymous):

1)|slope(AC)|=|slope(DC)| 2)AB^2+BC^2=AC^2

OpenStudy (anonymous):

alternatively for (2), to check that the angle=90degrees, you can use the dot product (A-B).(B-C)=0

OpenStudy (anonymous):

solve equations 1,2 for a,b

OpenStudy (anonymous):

Why is |slope (AC)| = |slope(DC)|? Also, is No. 1 like this? \[|\sqrt{a^2 - 4a + b^2 - 2b + 5}| = |\sqrt{5}|\] I don't know how to do no. 1 after that.

OpenStudy (anonymous):

Since angle ABC = 90 degrees Hence AB must be Perpendicular to BC For perpendicular lines we know that: Slope of AB * Slope of BC = -1 i.e. (b-0)/[a-(-1)] *(0-1)/(-1-2) =-1 i.e. b/(a+1) *(-1)/(-3) =-1 i.e. b/(a+1) *(1)/(3) =-1 i.e. b/(a+1) =-3 i.e. b = -3(a+1)------(1) Again, since given triangle is right angled, hence by Pythagoras Theorem and distance formula we have; \[AC^2= AB^2+BC^2\] A(a, b), B (-1,0), C (2, 1) \[(a-2)^2+(b-1)^2= (a+1)^2+(b-0)^2+(-1-2)^2+(0-1)^2\] \[a^2-4a+4+b^2-2b+1= a^2+2a+1+b^2+9+1\] \[-4a+4-2b+1= 2a+1+9+1\] \[-4a-2b+5= 2a+11\] \[-2b= 4a+2a+11-5\] \[-2b= 6a+6\] \[b= -3a-3\] \[b= -3(a+1)\]-----(2 From equations (1 and (2 we have